# Trickle-down Theorems via C-Lorentzian Polynomials II: Pairwise Spectral Influence and Improved Dobrushin’s Condition

### v2

@@ Abstract
We show that if the largest eigenvalue of the pairwise spectral influence matrix with entries ${\cal I}_{u,v}$ is bounded away from 1, i.e. $\lambda_{\max}({{\cal I}})\leq 1-\epsilon$ (and $X$ is connected), then the Glauber dynamics mixes rapidly and generate samples from $\mu$. This improves/generalizes the classical Dobrushin’s influence matrix as the ${\cal I}_{u,v}$ lower-bounds the classical influence of $u\to v$. As a by-product, we also prove improved/almost optimal trickle-down theorems for partite simplicial complexes. The proof builds on the trickle-down theorems via $\cal C$}-Lorentzian polynomials machinery recently developed by the authors and Lindberg.
@@ 1.1 Dobrushin’s Condition
Spectral independence is recently introduced in [ALO24, Che+21, Fen+21] and it was shown that it implies that the Glauber dynamics mixes in $O(d^{1/\eta})$-steps.
@@ 1.2 Main Results
For any connected $d$-partite complex, if $\lambda_{\max}({\cal I})\leq 1-\epsilon$ then for any $\tau\in X$ of codimension at least 2, we have $\lambda_{2}(P_{\tau})\leq\frac{(1-\epsilon)^{2}}{(\operatorname{codim}(\tau)-1)\epsilon}$. That is, the distribution $\mu$ is $\frac{(1-\epsilon)^{2}}{\epsilon}$-spectrally independent.
@@
One can then use this with Theorem 1.4 to prove that $\lambda_{2}(P_{\tau})\leq\frac{\lambda_{max}(M_{\tau})}{d-1}=\frac{1}{2}$ for all $\tau\in X$, which is one of the main results of [LLO25]. We give a more detailed proof of this in Section 5.3.
@@ 1.3 Applications
Example.
Consider the following concrete example: We are given a family of $d$ individuals (sites) on a graph $G=(V,E)$ where each edge represents mutual friendship. Suppose a spin represents the activity that an individual is doing (during the day) such as working, sleeping etc. A new contagious disease is started in the community; when an individual gets sick, we mark them as “sick” and we will immediately send their friends to “quarantine”. The question is to sample a uniformly random state of such a system. That is for a $\sigma\in X$ let
$\mu(\sigma)\propto\prod_{u\sim v}{\mathbb{I}}[s_{u}\neq\text{``sick''}\vee s_{v}=\text{``quarantine''}]$
The following lemma is immediate:
Lemma 1.6.
Suppose $\Delta$ is the maximum degree of any individual in $G$ and suppose $|S_{v}|\geq q$ for every site $v$. There is a universal constant $c>0$ such that if $\Delta<c\sqrt{q}$, then $\lambda_{\max}({\cal I})\leq 1/2$ and the Glauber dynamcis mixes in polynomial time.
We emphasize that for any $u\sim v$ the influence of $v$ to $u$ is almost 1, $I_{v\to u}\geq 1-O(1/q).$ So, $\lambda_{\max}(I)$ is not bounded and the classical (path) coupling arguments/Dobrushin’s like conditions fails to analyze the Glauber dynamics in such settings.
@@ 1.5 Acknowledgements
The first author acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2023-03726]. Cette recherche a été partiellement financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), [numéro de référence RGPIN-2023-03726].
@@ 5.2 Generalizing Dobrushin’s Condition
${\mathbb{E}}_{Q\sim{\cal W}_{V(\sigma)}[u]}[|Q|-1]=\sum_{k=2}^{\infty}{\mathbb{P}}[|Q|\geq k]\leq\sum_{k=2}^{\infty}(1-\epsilon)^{k}=\frac{(1-\epsilon)^{2}}{\epsilon}$
where in the first identity we used that every walk $Q$ has length at least 1. The inequality uses that from every vertex $v\in V_{i}$ we jump to an absorbing vertex $b_{v}$ with probability at least $\epsilon_{i}\geq\epsilon$. This then implies $\lambda_{2}(P_{\sigma})\leq\frac{(1-\epsilon)^{2}}{(\operatorname{codim}(\sigma)-1)\epsilon}.$ Therefore $\mu$ is $\frac{(1-\epsilon)^{2}}{\epsilon}$-spectrally independent. ∎

### v3

@@ Abstract
We show that if the largest eigenvalue of the pairwise spectral influence matrix with entries ${\cal I}_{u,v}$ is bounded away from 1, i.e. $\lambda_{\max}({{\cal I}})\leq 1-\epsilon$ (and $X$ is connected), then the Glauber dynamics mixes rapidly and generate samples from $\mu$. This improves/generalizes the classical Dobrushin’s influence matrix as the ${\cal I}_{u,v}$ lower-bounds the classical influence of $u\to v$. As an application, we prove that the Glauber dynamics mixes rapidly up to (approximately) the phase transition for the multi-state hard core model – a widely studied model in telecommunication networks and statistical physics (generalizing the hard core model) introduced by Mazel and Suhov. As a by-product of our results, we also prove improved/almost optimal trickle-down theorems for partite simplicial complexes. Our proof builds on the trickle-down theorems via $\cal C$}-Lorentzian polynomials machinery recently developed by the authors and Lindberg.
@@ 1.1 Dobrushin’s Condition
Spectral independence is recently introduced in [ALO24, Che+21, Fen+21] and it was shown that it implies that the Glauber dynamics mixes in $O(d^{\eta+1})$-steps.
@@ 1.2 Main Results
For any connected $d$-partite complex, if $\lambda_{\max}({\cal I})\leq 1-\epsilon$ then for any $\tau\in X$ of codimension at least 2, we have $\lambda_{2}(P_{\tau})\leq\frac{1-\epsilon}{(\operatorname{codim}(\tau)-1)\epsilon}$. That is, the distribution $\mu$ is $\frac{1-\epsilon}{\epsilon}$-spectrally independent.
@@
One can then use this with Theorem 1.4 to prove that $\lambda_{2}(P_{\tau})\leq\frac{\lambda_{\max}(M_{\tau})}{d-1}=\frac{1}{2}$ for all $\tau\in X$, which is one of the main results of [LLO25]. We give a more detailed proof of this in Section 5.3.
@@ 1.3 Applications
The Multi-state Hard Core Model.
Given a graph $G=(V,E)$, and a parameter $\lambda>0$, in the hard core model, the probability of each independent set $I$ of $G$ is proportional to $\lambda^{|I|}$, Studying the hardcore model has been pivotal in helping to understand the relationship between phase transitions in statistical physics and phase transitions in efficient approximability, see e.g., [ALO24, CLV21] for recent applications of the spectral independence machinery in studying the hardcore model.
The multi-state hard core model is a multi-state spin system generalization of the hard core model [Lou91, MS91]. Given an additional parameter $C\in{\mathbb{Z}}_{>0}$, every vertex $v\in V$ is now associated with the set of spins $S_{v}=\{0,1,\ldots,C\}$, and a configuration of spins $\{s_{v}\}_{v\in V}$ is admissible if for every edge $\{u,v\}\in E$ we have $s_{u}+s_{v}\leq C$. The probability of such a configuration is proportional to $\lambda^{\sum_{v}s_{v}}$. Observe that when $C=1$ this model coincides with the hard core model, where the vertices with spin $s_{v}=1$ give an independent set of $G$. The model was first introduced in [Lou91] to study telecommunication networks. Later, it was observed by Mazel and Suhov [MS91] that the model can be treated as a ”subensemble” of the well-known Solid-on-Solid (SOS) model in statistical physics. Here we omit the details and refer interested readers to [Cap+14] for background information on the SOS model and connections to statistical physics and sampling.
Galvin, Martinelli, Ramanan and Tetali [Gal+11] observed that the model undergoes a phase transition at $\lambda\approx\left(\frac{\log\Delta}{(C+2)\Delta}\right)^{2/(C+2)}$ for even $C$ and at $\lambda\approx\left(\frac{e}{\Delta}\right)^{2/(C+1)}$ for odd $C$, when $G$ is a $\Delta$-ary tree. To the best of our knowledge, despite significant applications, no prior counting and sampling results are known for the multi-state hard core model. Here we use Theorem 1.3 to efficiently sample a random configuration of the multi-state hard core model for $\lambda$ up to almost the uniqueness threshold.
Theorem 1.6.
Given a instance of the multi-state hard core model, i.e., a graph $G=(V,E)$, $0\leq\lambda<1$ and $C\in\mathbb{Z}_{>0}$, then for any $u,v\in V$,
${\cal I}(u,v)\leq\frac{\lambda^{\frac{C+1}{2}}}{1-\sqrt{\lambda}}.$
Therefore, if $\lambda_{\max}({\cal I})\leq 1-\epsilon$, $\mu$ is $\frac{1-\epsilon}{\epsilon}$-spectrally independent and the Glauber dynamics mixes rapidly.
A few remarks are in order. If $\Delta$ is the maximum degree of $G$, then
  • It follows that the Glauber dynamics mixes rapidly as long as
    $\lambda\leq\left(\frac{1-\epsilon}{(C+1)\Delta}\right)^{2/(C+1)}.$
    Note that this converges to $1$ as $C\to\infty$ (and $G$ is fixed). This almost matches the phase transition for even $C$ as $C\to\infty$.
  • If we fix $\epsilon$ and $C$, then for $\Delta\geq\left(\frac{1}{\epsilon}\right)^{1/(C+1)}$ the Glauber dynamics mixes rapidly as long as
    $\lambda\leq\left(\frac{1-2\epsilon}{\Delta}\right)^{2/(C+1)}.$
    This almost matches the phase transition for odd $C$ as $\Delta\to\infty$.
  • The (classical) influence of $v$ on $u$ can be as large as $\lambda$ (in the worst case). So, the classical Dobrushin’s/coupling type arguments one can only prove fast mixing if $\lambda<\frac{1-\epsilon}{\Delta}$ (independent of $C$).
Proof.
Fix two sites $u,v\in V$, and let $\tau\in X(d-2)$ be any allowed choice of spins for all other sites except $u,v$. Then based on the spin of sites adjacent to $u,v$, there exist $C_{u},C_{v}\leq C$ such that $\tau\cup\{s_{u},s_{v}\}\in X(d)$ if and only if $s_{u}\leq C_{u}$, $s_{v}\leq C_{v}$, and $s_{u}+s_{v}\leq C$. With this, the probability transition matrix for the remaining vertices $u,v$ is $P_{\tau}=D^{-1}\left[\begin{smallmatrix}0&A\\ A^{\top}&0\end{smallmatrix}\right]$, where $A$ is the $C_{u}\times C_{v}$ matrix given by
$a_{s_{u}s_{v}}=\begin{cases}\lambda^{s_{u}+s_{v}},&s_{u}+s_{v}\leq C,\\ 0,&\text{otherwise}\end{cases}$
and $D$ is the diagonal matrix consisting of the row sums of $A$ and $A^{\top}$. By Lemma B.1, we have $\lambda_{2}(P_{\tau})\leq\lambda^{(C+1)/2}/(1-\sqrt{\lambda})$ as desired. ∎
@@ 1.5 Acknowledgements
We would like to thank Prasad Tetali for pointing us to possible applications of our results to the multi-state hard core model. The first author acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2023-03726]. Cette recherche a été partiellement financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), [numéro de référence RGPIN-2023-03726].
@@ 5.2 Generalizing Dobrushin’s Condition
${\mathbb{E}}_{Q\sim{\cal W}_{V(\sigma)}[u]}[|Q|-1]=\sum_{k=2}^{\infty}{\mathbb{P}}[|Q|\geq k]\leq\sum_{k=1}^{\infty}(1-\epsilon)^{k}=\frac{1-\epsilon}{\epsilon}$
where in the first identity we used that every walk $Q$ has length at least 1. The inequality uses that from every vertex $v\in V_{i}$ we jump to an absorbing vertex $b_{v}$ with probability at least $\epsilon_{i}\geq\epsilon$. This then implies $\lambda_{2}(P_{\sigma})\leq\frac{1-\epsilon}{(\operatorname{codim}(\sigma)-1)\epsilon}.$ Therefore $\mu$ is $\frac{1-\epsilon}{\epsilon}$-spectrally independent. ∎
@@ Appendix A Pairwise Influence vs the Spectral Pairwise Influence Matrices
Appendix B Sum Graphs
Let $G=(L,R,E)$ be a bipartite graph with $L=\{0,1,\dots,n_{L}-1\},R=\{n_{L},n_{L}+1,\dots,n_{L}+n_{R}-1\}$ for $n_{L},n_{R}\leq C+1$. Let $w:E\to{\mathbb{R}}$ defined as follows: For $i\in L,j\in R$ with $i+j-n_{L}\leq C$ there is an edge with weight $w(i,j)=\lambda^{i+j-n_{L}}.$
Lemma B.1.
If $\lambda<1$, then $\lambda_{2}(P_{G})\leq\frac{\lambda^{(C+1)/2}}{1-\sqrt{\lambda}}$, where $P_{G}$ is the transition probability matrix of the simple random walk on $G$
Proof.
Let $A$ be the adjacency matrix of $G$ and $D$ be the degree matrix. Note that,
$d_{w}(i)\geq\lambda^{i},~~\forall i\in L,\qquad\qquad d_{w}(j)\geq\lambda^{j-n_{L}},~~\forall j\in R.$ (9)
We need to show that
$A\preceq vv^{T}+\frac{\lambda^{(C+1)/2}}{1-\sqrt{\lambda}}D$ (10)
for some vector $v\in{\mathbb{R}}^{L+R}$.
Let $v_{L}\in{\mathbb{R}}^{L},v_{R}\in{\mathbb{R}}^{R}$ defined as follows: For any $i\in L$, $v_{L}(i)=\lambda^{i}$ and for any $j\in R$, $v_{R}(j)=\lambda^{j-n_{L}}$. Let
$B:=A-(v_{L}+v_{R})(v_{L}+v_{R})^{T}+v_{L}v_{L}^{T}+v_{R}v_{R}^{T}$
To prove (10), it is enough to show that
$B\preceq\frac{\lambda^{(C+1)/2}}{1-\sqrt{\lambda}}D.$ (11)
First, we observe that
$B=\sum_{i\in L,j\in R:i+j-n_{L}>C}B_{i,j},$
where $B_{i,j}$ is only supported on the $(i,j)$ and $(j,i)$ entries and it is the same as $B$ on those entries.
By Lemma B.2 below we can write
$B_{i,j}\preceq\lambda^{i+\frac{i+j-n_{L}}{2}}I_{i}+\lambda^{j-n_{L}+\frac{i+j-n_{L}}{2}}I_{j},$
where $I_{i}\in{\mathbb{R}}^{(L+R)\times(L+R)}$ is a diagonal matrix with $\mathbf{1}_{i}$ on the diagonal. Having this, we write,
$B\preceq\sum_{i,j:i+j-n_{L}>C}\left(\lambda^{i+\frac{i+j-n_{L}}{2}}I_{i}+\lambda^{j-n_{L}+\frac{i+j-n_{L}}{2}}I_{j}\right)\preceq\frac{\lambda^{(C+1)/2}}{1-\sqrt{\lambda}}D$
To see the last inequality, notice that for any $i\in L$ we have
$\sum_{j\in R:i+j-n_{L}>C}\lambda^{i+\frac{i+j-n_{L}}{2}}\leq\lambda^{i}\cdot\left(\lambda^{\frac{C+1}{2}}+\lambda^{\frac{C+2}{2}}+\dots\right)=\lambda^{i}\cdot\frac{\lambda^{(C+1)/2}}{1-\sqrt{\lambda}}$
where we used (9) in the last inequality. A similar inequality holds for any $j\in R$. This proves (11) as desired. ∎
Lemma B.2.
For any $\alpha\in{\mathbb{R}}$ and for any $a,b\geq 0$ such that $a\cdot b=\alpha^{2}$ we have
$\begin{bmatrix}0&\alpha\\ \alpha&0\end{bmatrix}\preceq\begin{bmatrix}a&0\\ 0&b\end{bmatrix}$
Proof.
The statement simply follows from when we subtract the left from the right we get a non-negative diagonal with a 0 determinant matrix rank 2 matrix. ∎